# Data Compression for Fast Online Stochastic Optimization

### v1

@@ A.3 Proofs of Section 5 results
A.4 Online clustering algorithm
We give details for the online clustering algorithm, introduced in Section 6. We keep track of two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. For simplicity, below we assume the number of microclusters to be $Q$ and the number of macroclusters to be $K$; if the number of data-points is smaller, i.e., $Q_{t}\leq Q$ and $K_{t}\leq K$, the arrival of new data-points result in new clusters.
@@ A.4 Online clustering algorithm
We initialize the problem by solving approximately the $k$-centers problem (10) with respect to $\hat{{\mathbf{P}}}^{0}$, allowing up to $Q$ clusters. This set of centers is denoted $A^{Q}_{0}$. We assign all points to the closest center, and define the support of the $q$-th cluster $S^{q}_{0}$ as the Voronoi region around its center $a_{0}^{q}\in A^{Q}_{0}$. Note that the Voronoi regions need not be explicitly stored; they are implicitly defined by their centers. For each cluster, we note its center, centroid, root-mean-squared-error (RMSE), and weight. These are denoted the microclusters. We refer to the RMSE of each cluster as $\eta^{q}_{t}$, and for clusters with only a single data-point, heuristically initialize it as twice the minimum RMSE of all clusters.
To create $K$ macroclusters, we solve approximately the $k$-centers problem with respect to $A^{Q}_{0}$, and obtain a set of centers $A^{K}_{0}$. Each microcluster is assigned to the macrocluster with the closest center; each macrocluster then contains all the points of its constituent microclusters, and has their combined weight. The centroid of the macrocluster is therefore the weighted average of the centroids of the constituent microclusters. The support of the $k$-th macrocluster is the Voronoi region around the $k$-th center $a_{0}^{k}\in A^{K}_{0}$, plus the data-points assigned to it, and minus the data-points assigned to other clusters, i.e.,
@@
In this manner, the set of supports $\mathcal{S}_{t}$ satisfy Assumption 3.2. The necessity of this adjustment of the support follows from the two-layer nature of the clustering algorithm. It is possible for a data-point to be clustered into the $q$-th microcluster, which is assigned to the $k$-th macrocluster, but distance-wise is actually closer to the center of $k^{\prime}$-th macrocluster for some $k^{\prime}\neq k$.
@@
For $t\geq 1$, when we observe a new data-point $\hat{u}$, we calculate its distance to the microcluster whose support it falls on, i.e., if it falls on support $S^{q}_{t-1},$
@@
If this value is below $2\eta^{q}_{t-1}$, where $\eta^{q}_{t-1}$ is the RMSE, we assign it to this cluster. The multiplier of 2 is selected according to CluStream sensitivity analysis [2, Section 6.4]. We then update all microclusters weights, as well as the centroid and RMSE of the assigned microcluster. The macroclusters are also adjusted accordingly, based on their constituent microclusters.
On the other hand, if $d_{t}>2\eta^{q}_{t-1}$, we create a new microcluster with this data-point as its center, and assign it the weight $1/n_{t}$. The weights of the other microclusters are decreased accordingly, following (17). The RMSE of this cluster is initialized heuristically to be twice the minimum RMSE of all other existing clusters. If the total number of microclusters exceeds $Q$ with this addition, we merge the two closest microclusters based on the distances between their centers. The center of the merged microcluster will be the weighted average of the centers of the two constituent microclusters, with the new centroid, RMSE, and weight calculated accordingly. The supports of the microclusters will be reassigned as in (20). Then, we again perform approximate $k$-centers on the $Q$ microclusters, warm-staring at the previous centers, and generate $K$ macroclusters. The parameters of the macroclusters are assigned the same manner as in initialization.
@@
In the algorithm and description, the cluster assignments are assumed to use the cluster centers. The centroid can also be used, however, up to time $t\leq\tau$.
@@
2:$A^{Q}_{0}\leftarrow$ find $Q$ centers using approximate $k$-centers on $\mathcal{D}_{0}$
3:assign all data-points to microclusters $C^{q}_{0}$, and compute $n^{q}_{0}$, $\theta^{q}_{0}$, $\bar{u}_{0}^{q}$, and $\eta^{q}_{0}$
4:$\hat{S}^{Q}_{0}\leftarrow$ (20)
5:$A^{K}_{0}\leftarrow$ find $K$ centers using approximate $k$-centers on $A^{Q}_{0}$
@@
10:     observe data-point $\hat{u}$, and set $d_{t}\leftarrow\|\hat{u}-a^{q^{\prime}}_{t-1}\|$, where $q^{\prime}$ is chosen such that $\hat{u}\in S^{q^{\prime}}_{t-1}$
@@
12:         $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ following  (16)
@@
14:         $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
@@
18:         $\bar{u}^{q^{\prime}}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\bar{u}^{q^{\prime}}_{t-1}+\hat{u})/n^{q^{\prime}}_{t}$$\triangleright$ update centroid

### v2

@@ A.2 Proofs of Section 4 results
A.3 Online clustering algorithm (Section 5.3)
We give details for the online clustering algorithm, introduced in Section 5.3. We keep track of two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. For simplicity, below we assume the number of microclusters to be $Q$ and the number of macroclusters to be $K$; if the number of datapoints is smaller, i.e., $Q_{t}\leq Q$ and $K_{t}\leq K$, the arrival of new datapoints result in new clusters.
@@
We initialize the problem by solving approximately the $k$-centers problem (11) with respect to $\hat{{\mathbf{P}}}^{0}$, allowing up to $Q$ clusters. This set of centers is denoted $A^{Q}_{0}$. We assign all points to the closest center, and define the support of the $q$-th cluster $S^{q}_{0}$ as the Voronoi region around its center $a_{0}^{q}\in A^{Q}_{0}$. Note that the Voronoi regions need not be explicitly stored; they are implicitly defined by their centers. For each cluster, we note its center, mean, root-mean-squared-error (RMSE), and weight. These are denoted the microclusters. We refer to the RMSE of each cluster as $\eta^{q}_{t}$, and for clusters with only a single datapoint, heuristically initialize it as twice the minimum RMSE of all clusters.
To create $K$ macroclusters, we solve approximately the $k$-centers problem with respect to $A^{Q}_{0}$, and obtain a set of centers $A^{K}_{0}$. Each microcluster is assigned to the macrocluster with the closest center; each macrocluster then contains all the points of its constituent microclusters, and has their combined weight. The mean of the macrocluster is therefore the weighted average of the means of the constituent microclusters. The support of the $k$-th macrocluster is the Voronoi region around the $k$-th center $a_{0}^{k}\in A^{K}_{0}$, plus the datapoints assigned to it, and minus the datapoints assigned to other clusters, i.e.,
@@
In this manner, the set of supports $\mathcal{S}_{t}$ satisfy Assumption 3.1. The necessity of this adjustment of the support follows from the two-layer nature of the clustering algorithm. It is possible for a datapoint to be clustered into the $q$-th microcluster, which is assigned to the $k$-th macrocluster, but distance-wise is actually closer to the center of $k^{\prime}$-th macrocluster for some $k^{\prime}\neq k$.
@@
For $t\geq 1$, when we observe a new datapoint $\hat{u}$, we calculate its distance to the microcluster whose support it falls on, i.e., if it falls on support $S^{q}_{t-1},$
@@
If this value is below $2\eta^{q}_{t-1}$, where $\eta^{q}_{t-1}$ is the RMSE, we assign it to this cluster. The multiplier of 2 is selected according to CluStream sensitivity analysis [2, Section 6.4]. We then update all microclusters weights, as well as the mean and RMSE of the assigned microcluster. The macroclusters are also adjusted accordingly, based on their constituent microclusters.
On the other hand, if $d_{t}>2\eta^{q}_{t-1}$, we create a new microcluster with this datapoint as its center, and assign it the weight $1/n_{t}$. The weights of the other microclusters are decreased accordingly, following (19). The RMSE of this cluster is initialized heuristically to be twice the minimum RMSE of all other existing clusters. If the total number of microclusters exceeds $Q$ with this addition, we merge the two closest microclusters based on the distances between their centers. The center of the merged microcluster will be the weighted average of the centers of the two constituent microclusters, with the new mean, RMSE, and weight calculated accordingly. The supports of the microclusters will be reassigned as in (22). Then, we again perform approximate $k$-centers on the $Q$ microclusters, warm-staring at the previous centers, and generate $K$ macroclusters. The parameters of the macroclusters are assigned the same manner as in initialization.
@@
In the algorithm and description, the cluster assignments are assumed to use the cluster centers. The mean can also be used, however, up to time $t\leq\tau$.
@@
[…]
3:assign all datapoints to microclusters $C^{q}_{0}$, and compute $n^{q}_{0}$, $\theta^{q}_{0}$, $\bar{u}_{0}^{q}$, and $\eta^{q}_{0}$
[…]
@@
10:  observe datapoint $\hat{u}$, and set $d_{t}\leftarrow\|\hat{u}-a^{q^{\prime}}_{t-1}\|$, where $q^{\prime}$ is chosen such that $\hat{u}\in S^{q^{\prime}}_{t-1}$
@@
12:   $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ following  (18)
@@
14:   $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
@@
18:   $\bar{u}^{q^{\prime}}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\bar{u}^{q^{\prime}}_{t-1}+\hat{u})/n^{q^{\prime}}_{t}\triangleright$ update mean